# Sensory Glove-Based Surgical Robot User Interface

### v1

@@ Abstract
Robotic surgery has reached a high level of maturity and has become an integral part of standard surgical care. However, existing surgeon consoles are bulky and take up valuable space in the operating room, present challenges for surgical team coordination, and their proprietary nature makes it difficult to take advantage of recent technological advances, especially in virtual and augmented reality. One potential area for further improvement is the integration of modern sensory gloves into robotic platforms, allowing surgeons to control robotic arms directly with their hand movements intuitively. We propose one such system that combines an HTC Vive tracker, a Manus Meta Prime 3 XR sensory glove, and God Vision wireless smart glasses. The system controls one arm of a da Vinci surgical robot. In addition to moving the arm, the surgeon can use fingers to control the end-effector of the surgical instrument. Hand gestures are used to implement clutching and similar functions. In particular, we introduce clutching of the instrument orientation, a functionality not available in the da Vinci system. The vibrotactile elements of the glove are used to provide feedback to the user when gesture commands are invoked. A preliminary evaluation of the system shows that it has excellent tracking accuracy and allows surgeons to efficiently perform common surgical training tasks with minimal practice with the new interface; this suggests that the interface is highly intuitive. The proposed system is inexpensive, allows rapid prototyping, and opens opportunities for further innovations in the design of surgical robot interfaces.
@@ I INTRODUCTION
Robotic surgery has reached a high level of maturity and has become an integral part of standard surgical care. For the surgeon, it offers increased precision and dexterity through efficient interaction interfaces. However, despite its high acceptance and success, it faces some challenges in modern operating rooms. The surgeon consoles are bulky and take up valuable space [1]. A redesigned workspace promises to improve turnaround time, increase throughput, and reduce cost [2]. Studies further emphasize the potential of ergonomic workspace design and nonoperative process redesign to advance the efficiency of robotic surgical devices [3, 4]. Another challenge with current consoles is that the surgeon is separated from the rest of the team, which can lead to challenges in communication [5] and can impact the efficiency of the operational workflow in tasks such as robotic docking [6]. Finally, the proprietary nature of existing consoles makes it difficult to innovate in this space despite rapid advances in fields such as augmented and virtual reality that could benefit the surgeon.
Several studies have attempted to address some of these shortcomings. [7] proposes a novel method for controlling the endoscope through a head-mounted master interface (HMI) integrated with a surgical robot system. [8] controls the surgical robot arm and the Augmented Reality (AR) interface through various hand gestures. In [5], the authors provide a nice review of existing approaches and opportunities.
One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allows surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles. Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
@@
In our work, we develop a system in which an HTC Vive Tracker [12] is attached to a Manus Meta Prime 3 Haptic XR sensory glove [13], equipped with multiple IMUs and bending sensors for finger tracking, to control one arm of the da Vinci surgical robot paired with the da Vinci Research Kit (dVRK) [14]; the visual feedback to the surgeon is provided through God Vision wireless smart glasses [15]. In addition, this setup manipulates the robotic arm’s end effector (jaw of the instrument) through finger movements. Moreover, we implemented hand gesture recognition to substitute pedal inputs of the current da Vinci console (e.g. clutch). This also allows us to clutch the instrument orientation, a functionality not provided by the da Vinci system. The glove is equipped with vibrotactile elements to provide feedback to the user about whether the gesture is activated.
We evaluate the proposed system in terms of tracking accuracy and the ability of the user to perform surgical training tasks similar to those used in training with the da Vinci devices. The evaluation shows that the system is responsive and that surgeons can effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive. The proposed system is inexpensive as it uses off-the-shelf components and it dramatically reduces space requirements. It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
@@ II METHODOLOGY
Figure 1: The setup showing the configuration of the da Vinci arm (PSM) and the Manus glove with the HTC Vive Tracker attached. The user wears the God Vision smart glasses for the stereo endoscopic view. Shown are the $x$ (red), $y$ (green), and $z$ (blue) axes of the relevant coordinate frames.
@@ II-A System Architecture
The system integrates several components. The surgical robot platform consists of the first-generation da Vinci robot paired with the dVRK. One of the Patient Side Manipulators (PSMs) is equipped with the Fenestrated Bipolar Forceps (FBF) instrument and is subject to glove control. The arm is configured as in Fig. 1. The HTC Vive Tracker attached to the Manus glove provides the pose of the hand, while the fingers are tracked by the Manus Meta Prime 3 XR. In addition, classifier models were trained for hand gesture recognition; hand gestures were used to supersede da Vinci’s pedal inputs. The visual feedback to the user is provided through God Vision smart glasses. The system runs the Robot Operating System (ROS) [16] that facilitates seamless communication and coordination between the Manus glove, the HTC Vive Tracker, and dVRK, ensuring real-time data streaming and responsive interaction between the user and the control interface. The architecture allows for easy prototyping.
@@
For the full da Vinci system, the kinematic chain from the Patient Side Manipulator (PSM) tip to the Endoscope Camera Manipulator (ECM) tip involves Setup Joints (SUJs). The poses provided by the dVRK exhibit errors within a range of ±$\pm 5cm$ for translation and fall between $5\sim 10$ degrees for orientation These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [14]. Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [17] that eliminates the influence of SUJs and allows us to accurately determine the transformation between the PSM base (frame $R$) and the instrument tip (frame $T$) (Fig. 1).
II-C Glove Kinematics
To monitor the glove’s position and orientation within the environment, we attached an HTC Vive Tracker to it. The glove’s position and orientation data are updated at 120 Hz using the data from the IMUs and HTC Vive Base Stations [18, 19].
The coordinate frames for the HTC Vive Tracker and Base Station are depicted in Fig. 1. At initialization, the tracker frame $H$ and the tracker base frame $S$ coincide and are aligned with the PSM instrument tip frame $T$. A displacement of $H$ from $S$ results in the scaled motion of PSM (see below) to create an intuitive control for the user.
II-D Control Architecture
To ensure safety for both the glove and the instrument tip, we define their operational workspaces to be cubes, with a larger side $L_{h}$ for the glove and a side $L_{t}$ for the instrument tip, centered at their initial positions. In both cases, movements beyond the cube volume are projected to the cube surface.
The movements of the glove tend to be significantly larger than the desired movements of the PSM tip ($L_{h}>L_{t}$). The motion of the glove is thus scaled to match the range needed for PSM operations:
@@ II-D Control Architecture
where $\Delta p_{h}$ represents the translational movement of the glove from its initial position, and $\Delta p_{t}$ denotes the expected translational movement of the PSM from its initial position, and $\eta={L_{t}}/{L_{h}}$ is the scaling factor. Similarly to the existing surgeon console, the scaling factor $\eta$ can be adjusted.
II-E Finger Tracking
@@
The tracking of the relative transformation between the fingers and the wrist is accomplished using the five sets of flexible sensors with 2 Degrees of Freedom (DoF) and six sets of 9 DoF IMUs of the Manus Meta Prime 3 Haptic XR glove, operating at a rate of 90Hz. These sensors boast a ±$\pm 2.5^{\circ}$ accuracy [13]. The glove software provides a hand skeleton mesh, inspired by the Mediapipe framework [20]. This mesh is dynamically adjusted based on the readings from the glove sensors, allowing for a real-time representation of the hand pose.
In our architecture, the Euclidean distance between the tips of the index and thumb fingers is used to control the jaw of the end effector. To map the distance of the fingers to the distance between the tips of the jaw we use an expression analogous to Eq. (1). This enables precise and intuitive control of the robotic system.
II-F Hand Gesture Recognition
@@
Hand gestures serve as a seamless and intuitive means of interaction and open up the possibility of introducing a large variety of inputs that can surpass what the current consoles can provide. For example, in [8] it was shown that it is possible to implement 8 different commands.
@@
In our work, we train 4 different multiclassification models for hand gesture recognition using our custom dataset. Our implemented architecture can currently distinguish between five distinct gestures: None, Pinky (thumb and pinky in contact), Ring (ring finger and thumb in contact), Fist, and Thumb up, as illustrated in Table I. Our framework allows a new gesture to be added in just a few minutes. Currently, only three of the five gestures serve as commands (Fist, Ring, and Pinky). Fist activates the clutch mechanism, Ring, if held for 2 seconds, initiates and suspends robot tracking, and Pinky triggers monopolar current delivery [21]. See Section III for further details.
II-G Clutch
@@
The clutch functionality in the da Vinci surgical robot serves as a safety feature and a means of control for the surgeon during robotic-assisted surgeries. This mechanism essentially allows the surgeon to recenter the workspace of the surgeon console by temporarily disengaging the movement of the robotic arms while still maintaining control of the Master Tool Manipulators.
When the surgeon operates the da Vinci system’s master console, which consists of hand and finger controls, they can engage or disengage the clutch as needed. This allows them to stop the movement of the robotic arms, enabling them to make adjustments or reposition the controllers as required during the procedure.
@@
In the dVRK system, the clutch is engaged with a foot pedal or a button accessible by the index finger. However, the current rotation of the tool is maintained and cannot be changed (this is achieved by blocking the last two joints of the controller). In contrast, our system implements full clutching (both translation and rotation). The clutch is engaged by closing the hand (Fist gesture) as described next.
@@
Leading Edge: When the Fist gesture is detected, the clutch is engaged, and the vibrotactile feedback of the glove is activated. The current configuration of the PSM $g_{rt}^{0}$ is recorded and the robot is prevented from moving. Note that the jaw of the instrument is allowed to move.
Active Clutch: The vibrotactile feedback continues to be activated. The PSM is held immobile. The glove can be moved to an arbitrary pose, allowing the user to move to a comfortable configuration.
Trailing Edge: When the gesture becomes other than Fist, the clutch is disengaged and the vibrotactile feedback is deactivated. The tracker frame $H$ and the tracker base frame $S$ are aligned, and their axes are aligned with the axes of the PSM tip frame $T$ whose configuration is described by $g_{rt}^{0}$.
@@ III Hand Gesture Recognition
Hand Gesture Representation Hand Gesture Representation
Thumbs up [Uncaptioned image] Pinky [Uncaptioned image]
Ring [Uncaptioned image] Fist [Uncaptioned image]
TABLE I: Gestures trained for distinct functionalities. In this paper, Fist invokes the clutch mechanism and Ring toggles the robot control on or off. None gesture corresponds to the case when the subject is not using any of the four gestures above.
To train the classifiers for hand gesture recognition, we created a custom dataset. We used 5 sets of approximately $3-5$ minute recordings from the glove worn by one of the authors while performing one of the 4 gestures reported in Table I, or none of them. The recorded data (Table II) consists of the point cloud of each of the 21 landmark relative poses (3 position entries and 4 quaternion entries) regarding the wrist IMU (147 features). Table III shows examples of point clouds for different gestures. Each recording has been made while changing the position and orientation of the hand as much as possible to ensure the widest variety of data. Each recording corresponds to one gesture (or no gesture) so when merging all the data it was easy to attribute a class to each sample. As the number of classes is small, a simple one-hot encoding has been used.
Hand Gesture None Pinky Ring Fist Thumbs up
Train set $20743$ $12833$ $25011$ $16738$ $19447$
Test set $4167$ $2493$ $5075$ $3445$ $3775$
TABLE II: The size of our training and testing sets for hand gesture recognition.
Input Hand Pose Predicted Gesture Input Hand Pose Predicted Gesture
[Uncaptioned image] Thumbs-up [Uncaptioned image] Pinky
[Uncaptioned image] Ring [Uncaptioned image] Fist
TABLE III: Examples of point clouds used for gesture prediction.
@@ III-A Classifiers
We initially used Random Forests (RF) [22] as a model for gesture recognition. Subsequently, we explored the Multilayer Perceptron (MLP) classifier [23]. The MLP model was configured with two hidden layers containing 40 and 25 neurons, respectively, and a maximum of 100 iterations. The use of such a small model with only two hidden layers helps prevent overfitting.
In addition to these established models, we investigated a novel classifier, the Light Gradient Boosting Machine (L-GBM) [24]. L-GBM employs a sequential approach in which decision trees are added successively, with each tree correcting errors from the ensemble of preceding trees. Unlike traditional gradient-boosting methods, L-GBM directly minimizes the loss function. Tailored for efficient training on large datasets, it is particularly suitable for applications with resource constraints, and it can quickly predict data.
We also trained a Histogram-based Gradient Boosting model that combines the traditional gradient boosting [25] and the histogram-based algorithm [24]. Unlike traditional gradient boosting methods that operate on individual data points, histogram-based gradient boosting works with binned feature values. By constructing histograms of input features, this approach accelerates the computation of information gained during tree construction. The binning process reduces the number of unique feature values, resulting in faster training times and decreased memory usage.
III-B Results
By training the models above, it is possible to recognize 5 different gestures (tested on a new, unseen dataset). The models were implemented in ROS and Table IV shows their performance. The confusion matrices for each model are shown in Fig. 2.

### v2

@@ Abstract
Robotic surgery has reached a high level of maturity and has become an integral part of standard surgical care. However, existing surgeon consoles are bulky, take up valuable space in the operating room, make surgical team coordination challenging, and their proprietary nature makes it difficult to take advantage of recent technological advances, especially in virtual and augmented reality. One potential area for further improvement is the integration of modern sensory gloves into robotic platforms, allowing surgeons to control robotic arms intuitively with their hand movements. We propose one such system that combines an HTC Vive tracker, a Manus Meta Prime 3 XR sensory glove, and SCOPEYE wireless smart glasses. The system controls one arm of a da Vinci surgical robot. In addition to moving the arm, the surgeon can use fingers to control the end-effector of the surgical instrument. Hand gestures are used to implement clutching and similar functions. In particular, we introduce clutching of the instrument orientation, a functionality unavailable in the da Vinci system. The vibrotactile elements of the glove are used to provide feedback to the user when gesture commands are invoked. A qualitative and quantitative evaluation has been conducted that compares the current device with the dVRK console. The system is shown to have excellent tracking accuracy, and the new interface allows surgeons to perform common surgical training tasks with minimal practice efficiently.
@@ I INTRODUCTION
Robotic surgery has reached a high level of maturity and has become an integral part of standard surgical care. For the surgeon, it offers increased precision and dexterity through efficient interaction interfaces. However, despite its high acceptance and success, it faces challenges in modern operating rooms (ORs). The surgeon consoles are bulky and take up valuable space [1]. A redesigned workspace promises to improve turnaround time, increase throughput, and reduce cost [2]. Studies further emphasize the potential of ergonomic workspace design and nonoperative process redesign to advance the efficiency of robotic surgical devices [3, 4]. Another challenge with current consoles is that the surgeon is separated from the rest of the team, which can lead to challenges in communication [5] and can impact the efficiency of the operational workflow in tasks such as robotic docking [6]. Finally, the proprietary nature of existing consoles makes it challenging to innovate in this space despite rapid advances in fields such as augmented and virtual reality that could benefit the surgeon.
Several studies have attempted to address some of these shortcomings. In [7], a novel method is proposed to control the endoscope through a head-mounted interface (HMI) integrated with a surgical robot system. And [8, 9] control the surgical robot arm and the Augmented Reality (AR) interface through various hand gestures.
One potential area for further improvement and technological innovation in robotic surgery is the integration of modern sensory gloves into robotic platforms, which allow surgeons to intuitively control robotic arms with their hand movements without the need for additional manipulators that are customary for existing consoles. Furthermore, most end-effectors of surgical robot instruments, such as the Fenestrated Bipolar Forceps (FBF) and the Large Needle Driver (LND), are limited to the use of opposing blades due to the design of the console manipulators (as the last joint is controlled only by pinching two fingers); glove-based control could offer the surgeon additional dexterity while using these instruments.
@@
In our work, we develop a system in which an HTC Vive Tracker [13] is attached to a Manus Meta Prime 3 Haptic XR sensory glove [14], equipped with multiple IMUs and bending sensors for finger tracking to control one arm of the da Vinci surgical robot paired with the da Vinci Research Kit (dVRK) [15]. Visual feedback to the surgeon is provided through SCOPEYE wireless smart glasses [16]. In addition, this setup manipulates the robotic arm’s end effector (jaw of the instrument) through finger movements. Moreover, we implemented hand gesture recognition to substitute pedal inputs of the current da Vinci console (e.g., clutch). One of the gestures is employed to clutch the instrument orientation, a functionality not provided by the da Vinci system. The glove has vibrotactile sensors to inform the user that a gesture has been activated.
We evaluated the proposed system by having five surgeons and two medical doctors (no completed residency in surgery) perform a modified version of a robot-assisted surgical exercise using our system and the dVRK console. The system has been evaluated in terms of tracking accuracy and movement efficiency. After completing the exercise, the participants completed the System Usability Scale [17] and the NASA Task Load Index [18] questionnaires to measure the usability and workload of our proposed architecture. These evaluations show that the system is responsive and that surgeons can effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive. The proposed system is inexpensive, as it uses off-the-shelf components, and it dramatically reduces space requirements. Our work also demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
@@ II METHODOLOGY
Figure 1: The setup showing the configuration of the da Vinci arm (PSM) and the Manus glove with the HTC Vive Tracker attached. The user wears the SCOPEYE smart glasses for the stereo endoscopic view. Shown are the $x$ (red), $y$ (green), and $z$ (blue) axes of the relevant coordinate frames.
@@ II-A System Architecture
The system integrates several components. The surgical robot platform comprises the first-generation da Vinci robot paired with the dVRK. One of the Patient Side Manipulators (PSMs) is equipped with the Fenestrated Bipolar Forceps (FBF) instrument and is subject to glove control. The arm is configured as in Fig. 1. The HTC Vive Tracker attached to the Manus Meta Prime 3 XR glove provides the pose of the hand, while the fingers are tracked by the glove. Instead of using image-based hand-tracking devices as in [9], we decided to use sensor gloves because the hand-tracking performance decreases in the presence of visual occlusion or changes in room lighting. These occur frequently in the OR, such as when other surgical team members use secondary screens to look at details in the scene. Moreover, sensory gloves offer vibrotactile feedback. In addition, classifier models were trained for hand gesture recognition; hand gestures were implemented to substitute the da Vinci console’s pedal inputs. SCOPEYE smart glasses that provide the same visual feedback as the High-Resolution Stereo Video (HRSV) of the da Vinci console and have a maximum latency of 7ms [16] were used to further simplify the setup.
The system runs the Robot Operating System (ROS) [19] that facilitates seamless communication and coordination between the Manus glove, the HTC Vive Tracker, and dVRK, ensuring real-time data streaming and responsive interaction between the user and the control interface.
@@
For the full da Vinci system, the kinematic chain from the Patient Side Manipulator (PSM) tip to the Endoscope Camera Manipulator (ECM) tip involves Setup Joints (SUJs). The poses provided by the dVRK exhibit errors within a range of ±$\pm 5cm$ for translation and fall between $5\sim 10$ degrees for orientation. These inaccuracies arise due to challenges in calibrating the SUJs precisely, as reported in [15]. Consequently, we implemented a custom calibration method for the dVRK using fiducial markers [20] that eliminates the influence of SUJs and allows us to accurately determine the transformation between the PSM base (frame $R$) and the instrument tip (frame $T$), see Fig. 1.
II-C Scaling Workspace
To ensure safety for both the glove and the instrument tip, we define their operational workspaces as cubes, with a larger side $L_{h}$ (30cm) for the glove and a side $L_{t}$ (7.5cm) for the instrument tip, centered at their initial positions. In both cases, movements beyond the cube volume are projected to the cube surface.
The glove movements tend to be significantly larger than the desired movements of the PSM tip ($L_{h}>L_{t}$). The motion of the glove is thus scaled to match the range needed for PSM operations:
@@ II-C Scaling Workspace
where $\Delta p_{h}$ represents the translational movement of the glove from its initial position, $\Delta p_{t}$ denotes the expected translational movement of the PSM from its initial position, and $\eta={L_{t}}/{L_{h}}$ is the scaling factor. Similarly to the existing surgeon console, the scaling factor $\eta$ can be adjusted.
II-D Glove Kinematics
To monitor the glove’s position and orientation within the environment, we attached a HTC Vive Tracker to it. The glove position and orientation data are updated at 120 Hz using the data from the tracker’s IMUs and HTC Vive Base Stations [21, 22].
@@ II-D Glove Kinematics
Tracking of the relative transformation between the fingers and the wrist is accomplished using the five sets of flexible sensors with 2 Degrees of Freedom (DoF) and six sets of 9 DoF IMUs of the Manus glove, operating at 90Hz with ±$\pm 2.5^{\circ}$ accuracy [14]. The glove software provides a hand skeleton mesh inspired by the Mediapipe framework [23]. This mesh is dynamically adjusted based on the readings from the glove sensors, allowing for a real-time representation of the hand pose.
Hand Gesture Representation Hand Gesture Representation
Ring [Uncaptioned image] Clutch [Uncaptioned image]
TABLE I: Gestures trained for distinct functionalities. Clutch invokes the clutch mechanism and Ring turns the robot control on or off. If neither of these two gestures is recognized, the system returns None.
The HTC Vive Tracker and Base Station coordinate frames are depicted in Fig. 1. At initialization, the tracker frame $H$ and the tracker base frame $S$ coincide and are aligned with the PSM instrument tip frame $T$. The desired translation for the PSM end-effector is calculated by scaling the displacement of $H$ from $S$ as described above to create an intuitive control for the user.
For the goal orientation of the PSM end effector, we determine the relative orientation between the metacarpophalangeal (MP) and distal interphalangeal (DIP) joint of the sensory glove’s index finger and add this rotation to the one provided by the HTC Vive Tracker readings. This ensures that the system follows the index finger of the user, making the system more intuitive and reducing user fatigue.
In our architecture, the Euclidean distance between the tips of the index and thumb fingers was used to control the jaw of the end effector. To map the distance of the fingers to the distance between the tips of the jaw, we use an expression analogous to Eq. (1). The maximum opening of the jaw is achieved when the distance between the two fingertips is $10$cm. This value has been selected as it is the $95\%$ percentile of the length of the index finger of women [24], ensuring that all users can precisely and intuitively control the instrument jaw within the full range of motion.
II-E Hand Gesture Recognition
@@
Hand gestures serve as a seamless and intuitive means of interaction and open up the possibility of introducing a wide variety of inputs surpassing what the current consoles can provide. For example, [8] shows that it is possible to implement 8 different commands.
@@
In our work, we train 4 different multiclassification models for hand gesture recognition using our custom dataset. Our implemented architecture can currently distinguish between three distinct gestures: None, Ring (ring finger and thumb in contact), and Clutch (middle, ring, and pinky fingers closed) as illustrated in Table I. Clutch activates the clutch mechanism described in Section II-F. Holding the Ring gesture for 2 seconds initiates/suspends robot tracking. This implementation is open to allow new gestures to be added with a relatively small dataset. Further details are provided in Section III.
II-F Clutch
Input Hand Pose Predicted Gesture Input Hand Pose Predicted Gesture
[Uncaptioned image] Ring [Uncaptioned image] Clutch
TABLE II: Examples of point clouds used for gesture prediction.
@@ II-F Clutch
The clutch functionality in the da Vinci surgical robot serves as a safety feature and a means of control for the surgeon during robotic-assisted surgeries. This mechanism essentially allows the surgeon to recenter the workspace of the surgeon console by temporarily disengaging the movement of the robotic arms while still maintaining control of the Master Tool Manipulators (MTMs).
@@
In the dVRK system, the clutch is engaged with a foot pedal. However, the current rotation of the tool is maintained and cannot be changed by fixing the orientation of the controllers. Our system implements full clutching (both translation and rotation). The clutch is engaged by performing the Clutch gesture as described next.
@@
  • Leading Edge: When the Clutch gesture is detected, the clutch is engaged, activating the vibrotactile feedback of the glove. The current configuration of the PSM $g_{rt}^{0}$ is recorded, and the robot is prevented from moving.
  • Active Clutch: The vibrotactile feedback continues, and the PSM is held steady. The user can change the pose of the glove to a comfortable configuration.
  • Trailing Edge: When the gesture changes, the clutch is disengaged, and the vibrotactile feedback is deactivated. The tracker frame $H$ and the tracker base frame $S$ are aligned to the axes of the PSM tip frame $T$, whose configuration is described by $g_{rt}^{0}$.
@@ III Hand Gesture Recognition
We created a custom dataset for training hand gesture recognition, recording approximately $3-5$ minutes of the glove worn by different users performing one of the gestures in Table I, including the None. The data consists of point clouds with 21 hand joint landmarks, capturing the relative orientations to the base, set when the glove is activated. The orientations are represented by quaternions with respect to the glove base, initialized when the glove is turned on, resulting in a total of 84 features. The translations were not added because they vary for the same hand posture with each launch. Using only rotations thus makes it easy to capture gestures of interest with only a few recordings. Table II shows examples of point clouds for different gestures. Each recording contains one gesture, and the pose of the hand is changed to be as diverse as possible. We added $150,000$ gestures with sets of random normalized quaternions to the None class to prevent the models from recognizing a random gesture as one of the gestures of interest (Ring or Fist). The final dataset consists of $15,557$ Fist instances, $17,036$ Ring instances, and $178,217$ None instances. A simple one-hot encoding was applied due to the small number of classes.
@@ III-A Classifiers
We initially used Random Forests (RF) [25] as a model for gesture recognition. Subsequently, we explored the Feedforward Neural Network (FNN) [26] classifiers analogous to those introduced in [27]. However, the networks have been specifically tuned for classification. The structure of the FNN consists of an input layer ($n_{0}$) sized according to the input characteristics, followed by two fully connected layers ($n_{1}$, $n_{2}$), where the number of neurons is proportional to the size of the previous layer ($n_{1}=2\cdot n_{0}$, $n_{2}=2/3\cdot n_{1}$). The final layer ($n_{f}$) has $3$ output to function as a classifier, predicting one of the three possible classes.
In addition to these established models, we investigated the Light Gradient Boosting Machine (L-GBM) [28].Unlike traditional gradient-boosting methods like XGBoost [29], where trees grow depth-wise, L-GBM trees grow leaf-wise, making training faster. Tailored for efficient training on large datasets, they are particularly suitable for applications with resource constraints and can quickly provide good predictions.
We also trained a Histogram-Based Gradient Boosting (HGBT) model that combines traditional gradient boosting [30] and a histogram-based algorithm [28]. The HGBT works with binned feature values rather than individual data points, which is the case with traditional gradient boosting methods. This approach accelerates the computation of information gained during tree construction by constructing histograms of input features. The binning process reduces the number of unique feature values, resulting in faster training times and decreased memory usage.
To optimize the model hyperparameters, we adopted Grid-Search Cross-Validation (GridSearchCV) [31]. First, we split the dataset into training and testing sets with the $8:2$ ratio. The search process used a 5-fold cross-validation strategy, in which each combination was evaluated across different data splits. The grid includes five hyperparameters with three different values, resulting in 1215 combinations after cross-validation. The models were evaluated on accuracy, and the best-performing combination of hyperparameters was selected.
III-B Training Results